OpenStudy (anonymous):

find f(a+1) x^2 + 2x +3

OpenStudy (anonymous):

\[if ~f \left( x \right)=x^2+2x+3,~then~f \left( a+1 \right)=\left( a+1 \right)^2+2\left( a+1 \right)+3\] solve it.